OpenStudy (anonymous):

Find the area of the region between the curves y=12-x^2 and y=x^2-6

OpenStudy (amistre64):

height times width = area right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

y1-y2 = distance between them, and dy = width

OpenStudy (anonymous):

But i dont know how to use it in this equations. im confused

OpenStudy (amistre64):

\[\int (12-x^2)-(x^2-6)dx\]

OpenStudy (amistre64):

i shoulda said dx to begin with but my mind was a bit off :)

OpenStudy (amistre64):

only thing missing is how far to go

OpenStudy (anonymous):

so integrate this and then what?

OpenStudy (amistre64):

then apply the limits that need to be calculated

OpenStudy (amistre64):

when does 12-x^2 =x^2-6 ??

OpenStudy (amistre64):

2x^2 = 18 x^2 = 9 x = +- 3

OpenStudy (anonymous):

so it'll that integral from -3 to 9

OpenStudy (amistre64):

\[\int_{-3}^{3} (12-x^2)-(x^2-6)dx\]or simply\[2\int_{0}^{3} (12-x^2)-(x^2-6)dx\]

OpenStudy (anonymous):

oh ok i understand ...thank you.

OpenStudy (anonymous):

so i just have to solve that integral?

OpenStudy (amistre64):

your welcome; and it might come ount negative, but that just means we subtracted in a mismatched order; ignore any signage in the answer and youll be fine

OpenStudy (amistre64):

the books always say: subtract the bottom curve from the top; but it really doenst matter in the end. One answer might be 9 while the other -9; so just ignore the sign

OpenStudy (anonymous):

thanks, do i do the same for Find the area of the region between the curves y=x^2 and y^2=x

OpenStudy (amistre64):

yes, but you would do best to get them speaking the same terms; sqrt(y) = x and y^2 = x for example

OpenStudy (amistre64):



OpenStudy (anonymous):

sqrt(y).. where did that come from?

OpenStudy (amistre64):

really? y = x^2 ; sqrt each side sqrt(y) = x

OpenStudy (anonymous):

lol oh, sorry.. totally forgot

OpenStudy (amistre64):

i know i know, you know how much algebra i had to forget to learn trig :)

OpenStudy (anonymous):

Lol, this stuff is totally confusing. so i would integrate x^2-x?

OpenStudy (amistre64):

that wouldnt be to practical; they are speaking different terms at the moment so are hard to reconcile. kinda like adding unlike denominators on fractions; we have to alter them to speak the same language

OpenStudy (amistre64):

y=x^2 and y^2=x y = x^2 y = sqrt(x) int (x^2 - sqrt(x)) from 0 to 1

OpenStudy (amistre64):

now that will turn out a negative answer; so ignore the sign